# SPAMming Labels: Efficient Annotations for the Trackers of Tomorrow

### v1

@@ 4.2 Implementation Details
Architecture and training. Our model consists of three main components: (i) a detector network, based on YOLOX [27], (ii) a ResNet-based ReID network [28], and (iii) a GNN hierarchy, all pretrained on the synthetic sequences of MOTSynth [23]. Please see the supplementary material for further details about training hyperparameters. After pretraining, we use the resulting model to obtain pseudo-labels for our target MOT dataset, and then fine-tune our detector and GNN on them. Note that our Model is never trained on human annotations, it is only trained on synthetic annotations and its own pseudo-labels.
Experimental setup. To assess the quality of SPAM’s labels, we use them to train two state-of-the-art online trackers: ByteTrack [78] and GHOST [57]. For validation, we follow the standard half-split setting [79, 78, 57] over MOT17 sequences. Annotation budgets are computed based on number of clicks as described in Section 3.3.
@@ 4.3 Synthetic Pretaining
Several works have shown that synthetic pre-training is a strong starting point for multi-object tracking methods [78, 23]. In this section, we aim to analyze the gap between training a model on synthetic data and training it with real data separately for the three most common tracking components: detection, association, re-identification. For each setting, only the specific component that we are analysing is trained on synthetic data while other components are all trained on real data. We perform this experiment on three datasets and plot the difference in HOTA when training on real data vs. synthetic data in Fig. 4 (a positive score indicates superior performance with real data training). Our first observation, is that the detector is the component that suffers the domain gap the most, with a performance drop of 9.9 HOTA points. Re-identification, on the other hand, has a similar performance when trained on synthetic or real data. Lastly, association performance is higher with real data training (2.1 HOTA). This first experiment gives us an important insight: labeling and retraining should be focused on detections and associations, and we can directly use the synthetically trained reID model.
@@ 4.4 Towards a Model for Labeling
Graph-based validation of detection candidates. In our framework, annotators have the capability to eliminate detections, but we refrain from requesting them to annotate new detections in order to maintain minimal annotation effort. Instead we start from an initial graph formed by an over-complete set of detections. Towards this end, we use a low confidence threshold for our detector, thereby reducing the number of false negatives in our initial graph and ensuring a high recall. As a result, this leads to an increase in the number of false positives incorporated into the graph, a tendency that is further accentuated by the fact that our detector is exclusively trained on synthetic data. We propose to leverage the graph structure itself to validate the detections based on temporal consistency. Specifically, we add an additional GNN-layer in our hierarchy, $GNN_{node}$, which intuitively examines the spatial and temporal consistency of the bounding box candidates and filters out implausible ones. In  Table 1, we assess the effect of this design choice. Model A uses only high confidence bounding boxes, model B creates a graph with also low confidence bounding boxes but without $GNN_{node}$, while C includes $GNN_{node}$. As we can see, simply adding low confidence boxes results in a drastic drop in performance on all metrics. Only by adding our $GNN_{node}$ we are able to take full advantage of the low confidence boxes without penalizing our performance with too many errors.
@@
SPAM is a SOTA tracker. Before moving to labeling with our Model, we want to verify its overall performance and generalization capabilities when trained on real data. Only by using a state-of-the-art model we will be able to reduce annotation effort as much as possible. We therefore pre-train our model on the synthetic dataset, and finetune it on the specific real dataset using groundtruth. For a fair comparison, we obtain object detections from a YOLOX detector [27] trained following [78, 57, 12]. Our model outperforms state-of-the-art by 1.0 HOTA point on MOT17, 1.1 HOTA points on MOT20, and 0.7 HOTA points on DanceTrack as shown in Table 2. Furthermore, on MOT17 and MOT20, there is also a significant improvement in IDF1, showing the identity-preserving capabilities of the model, which is especially important for our video annotation setting. Even if being state-of-the-art is not the main goal of this paper, having a strong performing model allows us to have an excellent starting point for producing pseudolabels. We would like to emphasize that, we use ground truth annotations only for this experiment just to validate the performance of our model with a fair comparison with other trackers. From this point onward, we will not use ground truth labels. We will use our synthetic pretraining weights and focus on the pseudolabeling and active learning aspects of our method to show how to annotate multi-object tracking datasets.
4.5 Enhancing Annotation Quality through Pseudolabels and Active Learning
Retraining our Model on Pseudolabels. In this experiment, we compare two models. The first one, SPAM, is only trained on synthetic data. The second one, SPAM , is trained in three stages: (i) pre-training on synthetic data, (ii) pseudo-labeling the real dataset with our model, and (iii) re-training our model on our own pseudo-labels. In  Section 4.5, we observe a significant improvement when using pseudo-labels to re-train our model: an increase of 4-6 HOTA, 4-6 MOTA and 4-7 IDF1 on MOT17, MOT20 and DanceTrack datasets. Note that these improvements come at zero human annotation effort as this process requires no manual annotations. Further pseudo-labeling the real dataset and re-training our model does not lead to another performance increase. It is therefore time to turn into manual annotations, and in particular, finding the best use of our annotator’s effort.
Model Pseudolabel HOTA $\uparrow$ MOTA $\uparrow$ IDF1 $\uparrow$
@@ 4.5 Enhancing Annotation Quality through Pseudolabels and Active Learning
Table 3: Performance boost obtained by our model when retraining with its own pseudolabels incurring no manual annotation cost.
@@
Active learning on graph hierarchies. We now test the effectiveness of our proposed active learning labeling on graph hierarchies, and show a comprehensive comparison with several baselines in Fig. 5. We directly measure the accuracy of the output labels given an increasing percentage of manual annotations selected by different acquisition functions on MOT17. The vertical axis shows the increase in HOTA over the baseline performance, i.e., the performance without active learning. Oracle is our upper bound and shows the performance if all possible labels are set to their correct value. Every method has two different labeling schemes: one for detections (nodes), and one for associations (edges) as explained in Section 3.3. In addition to our method, we explore active learning strategies commonly used in object detection literature, including Coreset [58], Ensemble [5], and Consistency over augmented inputs [21]. The naming convention denotes the active learning strategy for detection in the first part and the strategy for association in the second part. For instance, ’EntImg|Rand’ indicates image level entropy-based selection for detection and random selection for association.
The methods shown in red shades and circle markers compare three detection labeling methods while keeping association label selection random. Box-level labeling is far superior to commonly-used image-level labeling, or random labeling. This highlights the potential of performing active learning at the object level rather than the frame level. The methods in blue shades and triangle markers all use our proposed active learning selection on graph hierarchies for the association, and compare several detection selection schemes. The Coreset [58], which prioritizes diversity of the data over uncertainty, emerges as the poorest performing approach. This observation underlines that, when accurate predictions are available, an uncertainty-based approach proves to be a more effective complement to pseudo-labels. Box-level entropy, ensembles, and consistency all perform similarly. Finally, our method, with both detection and association label selection on graph hierarchies. Its performance is far superior to all compared baselines, staying at only 1.2 points from Oracle performance. Notably, we also unify the AL strategy for detection and association with our approach.
@@ 4.6 Putting Everything Together: SPAM Labels in Action
Model Labels Pseudolabels Active Learn. Ann. Effort $\downarrow$ HOTA $\uparrow$ MOTA $\uparrow$ IDF1 $\uparrow$
@@
Training trackers with our labels. In this experiment, we ablate the effects of our SPAM pipeline on the actual usecase of training a model with our provided labels. We ablate label quality under the following settings: (i) Synthetic training, (ii) Retraining our models on our Pseudo-labels, (iii) Active learning based acquisiton and annotation of a small percentage of real data. We train state-of-the-art online tracker ByteTrack [78] from scratch with the aforementioned sets of labels and report its performance on the MOT17 validation set in  Table 4. The goal performance we want to reach is the one achieve by the model trained on groundtruth annotations (row 1). Row 2 shows our model trained only on synthetic data, which is 3 HOTA points below our goal. Using pseudo-label based training of our model increases performance by almost a HOTA point, and the ground truth level performance is achieved by also using our active learning method to select which samples to label. By annotating only $3.3\%$ of the data, we can achieve the same results as the model trained on $100\%$ of the ground truth data. This shows the potential of our SPAM in reducing annotation effort for multi-object tracking datasets.
@@
How much data do we really need to label? In this final experiment, our goal is to assess the impact of our entire pipeline by answering the question: How much data do we actually need to label in order to achieve the same performance as a model trained on the entire groundtruth training dataset? For that, we choose two well-established trackers, ByteTrack [78] and GHOST [57] due to their online inference strategy and versatility, and train them from scratch using our SPAM pipeline. We report their performance on MOT17, MOT20, and DanceTrack test sets in  Table 5. We first confirm that our results on the validation set also generalize to the test set, and that we can reach GT performance with $3.3\%$ of annotation effort on MOT17. Note that for MOT17 our annotations yield a slight improvement over GT, which can be attributed to observed annotation noise in the original GT, particularly in moving sequences. Further details can be found in the supplementary material. MOT20 and DanceTrack are known to be more challenging datasets [60], which is confirmed by the fact that the trackers still miss a few points when trained on only $3.3\%$ of manually annotated data. Given that we cannot do a full analysis of different percentages because we are working on the test sets, we choose two percentages of manually annotated data on which to train our trackers: $10\%$ for MOT20, and $20\%$ for DanceTrack. In both cases, we achieve on-par performance with the trackers trained on full groundtruth data. This clearly shows the potential in properly leveraging synthetic data, pseudo-labels from a strong model, and active learning on graph hierarchies in order to label multi-object tracking data, where we can save $96.7\%$ of the annotation effort on MOT17, $90\%$ of the annotation effort on MOT20 and $80\%$ of the annotation effort on DanceTrack.
@@ 5 Conclusion
We have introduced SPAM, a labeling engine to efficiently obtain trajectory annotations from video. We have shown that SPAM is able to produce annotations that yield state-of-the-art tracking performance with a significantly reduced human annotation effort. Through our extensive experiments we have shown that (i) our synthetic pretraining + pseudo-label finetuning recipe yields a strong, annotation-free, labeler; (ii) our combination of active learning + pseudo-labeling for annotation generation yields high quality and efficient annotations; and (iii) that graphs are an excellent tool to provide comprehensive modelling for trajectory annotation. We believe that SPAM paves the way to make large-scale cheap annotation of tracking datasets a reality, which will in turn power the next generation of data-hungry tracking algorithms.
@@
Appendix A Additional Details about SPAM
A.1 GNN Model
Our hierarchical graph-neural network (GNN) utilizes the time-aware message passing framework introduced in [9]. We use a hierarchy of 10 GNNs operating over a video clip of 512 frames. At the first level, we employ a $GNN_{node}$, which considers a temporal window of 10 frames for each node. Subsequent levels, denoted as $GNN_{edge}$, cover 2, 4, …, up to 512 frames. We follow a simple and unified approach for all hierarchy levels: $GNN_{node}$ and $GNN_{edge}$ levels share the same input features and network architecture (as presented in Section 3.2 of the main paper). Moreover, $GNN_{edge}$ levels also share learnable weights across hierarchy levels. Overall, we rely on a lightweight GNN hierarchy with only approximately 65K parameters running at 20 frames per second (FPS).
A.2 Active Hierarchical Labeling
A.3 Implementation Details
Experimental setup. After obtaining labels with SPAM, we train two well-established online trackers ByteTrack [78] and GHOST [57] to evaluate our label quality in Section 4.6 of the main paper. Our main goal is to compare our label quality with ground truth label quality, therefore, we start trainings from scratch and only train on the dataset of interest (200 epochs for MOT17, 150 epochs for MOT20 and 100 epochs for DanceTrack). During both training and testing, we follow their default training and testing parameters [78, 57].
Appendix B Additional Experiments
B.1 Frame-based Labeling and FPS
@@
B.2 Label Noise for Moving Cameras
In our experiments detailed in Section 4.6 of the main paper, we found that on MOT17, our annotations with a budget of 3.3% yield slightly better performance compared to the ground truth. We attribute this observation to annotation noise in the original ground truth, particularly in sequences with moving cameras. To illustrate this, we provide detailed results on MOT17 for each sequence with ByteTrack and GHOST in Fig. 7. We compare relative HOTA scores between trainings on ground truth and SPAM labels. Positive scores indicate superior performance with SPAM labels, while negative scores indicate that ground truth labels result in better performance. Overall, with a 3.3% budget, ground truth exhibits better performance on static camera sequences, while SPAM outperforms ground truth labels on moving camera sequences. We attribute this to additional noise in ground truth labels on moving sequences caused by the interpolation-based labeling strategy reported in [18].

### v2

@@ 4.2 Implementation Details
Architecture and training. Our model consists of three main components: (i) a detector network, based on YOLOX [27], (ii) a ResNet-based ReID network [28], and (iii) a GNN hierarchy, all pretrained on the synthetic sequences of MOTSynth [23]. After pretraining, we use the resulting model to obtain pseudo-labels for our target MOT dataset, and then fine-tune our detector and GNN on them. Note that our Model is never trained on human annotations, it is only trained on synthetic annotations and its own pseudo-labels.
Experimental setup. Online trackers are real-time tools suitable for deployment on robots or autonomous vehicles, whereas offline trackers are better suited for labeling tasks. Therefore, we structure our experiments to utilize SPAM for labeling and use these labels with online trackers. To evaluate the quality of SPAM labels, we train two state-of-the-art online trackers with these labels: ByteTrack [77] and GHOST [56]. For validation, we follow the standard half-split setting [78, 77, 56] over MOT17 sequences. Annotation budgets are computed based on number of clicks as described in Sec. 3.3.
@@ 4.3 Synthetic Pretaining
Several works have shown that synthetic pre-training is a strong starting point for multi-object tracking methods [77, 23]. In this section, we aim to analyze the gap between training a model on synthetic data and training it with real data for the three most common tracking components: detection, association, re-identification. For each setting, only the specific component that we are analysing is trained on synthetic data while other components are all trained on real data. We perform this experiment on three datasets and plot the relative HOTA difference (a positive score indicates superior performance with real data training) when training on real data vs. synthetic data in Fig. 4. Our first observation, is that the detector is the component that suffers the domain gap the most, with a performance drop of 9.9 HOTA points. Re-identification, on the other hand, has a similar performance when trained on synthetic or real data. Lastly, association performance is 2.1 HOTA higher with real data training. This first experiment gives us an important insight: labeling and retraining should be focused on detection and association components, and we can directly use the synthetically trained reID model.
@@ 4.4 Towards a Model for Labeling
Graph-based validation of detection candidates. In our framework, annotators have the capability to eliminate detections, but we refrain from requesting them to annotate new detections in order to maintain minimal annotation effort. Instead we start from an initial graph formed by an over-complete set of detections. Towards this end, we use a low confidence threshold for our detector, thereby reducing the number of false negatives in our initial graph and ensuring a high recall. As a result, this leads to an increase in the number of false positives incorporated into the graph, a tendency that is further accentuated by the fact that our detector is exclusively trained on synthetic data. We propose to leverage the graph structure itself to validate the detections based on temporal consistency. Specifically, we add an additional GNN-layer in our hierarchy, $GNN_{node}$, which intuitively examines the spatial and temporal consistency of the bounding box candidates and filters out implausible ones. In  Tab. 1, we assess the effect of this design choice. Model A uses only high confidence bounding boxes, model B creates a graph with also low confidence bounding boxes but without $GNN_{node}$, while C includes $GNN_{node}$. As we can see, simply adding low confidence boxes results in a drastic drop in performance on all metrics. Only by adding our $GNN_{node}$ we are able to take full advantage of the low confidence boxes without penalizing our performance with too many errors.
@@
SPAM is a SOTA tracker. Before moving to labeling with our model, we want to verify its overall performance and generalization capabilities. Only by using a state-of-the-art model we will be able to reduce annotation effort as much as possible. We therefore pre-train our model on the synthetic dataset, and finetune it on the specific real dataset using groundtruth. For a fair comparison, we obtain object detections from a YOLOX detector [27] trained following [77, 56, 12]. Our model outperforms state-of-the-art by 1.0 HOTA point on MOT17, 1.5 HOTA points on MOT20, and 0.7 HOTA points on DanceTrack as shown in Tab. 2. Furthermore, on MOT17 and MOT20, there is also a significant improvement in IDF1, showing the identity-preserving capabilities of the model, which is especially important for our video annotation setting. Even if being state-of-the-art is not the main goal of this paper, having a strong performing model allows us to have an excellent starting point for producing pseudo-labels. We would like to emphasize that, we use ground truth annotations only for this experiment just to compare our model fairly with other trackers. From this point onward, we will not use ground truth labels. We will use our synthetic pretraining weights and focus on the pseudo-labeling and active learning aspects of our method to show how to annotate multi-object tracking datasets.
4.5 Enhancing Annotations with Pseudo-labels and Active Learning
Retraining our Model on Pseudo-labels. In this experiment, we compare two models. The first one, SPAM, is only trained on synthetic data. The second one, SPAM , is trained in three stages: (i) pre-training on synthetic data, (ii) pseudo-labeling the real dataset with our model, and (iii) re-training our model on pseudo-labels. In  Sec. 4.5, we observe a significant improvement when using pseudo-labels to re-train our model: an increase of 4-6 HOTA, 4-6 MOTA, and 4-7 IDF1 points on MOT17, MOT20, and DanceTrack, respectively. Note that these improvements come at zero human annotation effort as this process requires no manual annotations.
Model Pseudo-label HOTA $\uparrow$ MOTA $\uparrow$ IDF1 $\uparrow$
@@ 4.5 Enhancing Annotations with Pseudo-labels and Active Learning
Table 3: Performance boost obtained by our model when retraining with its own pseudo-labels incurring no manual annotation cost.
@@
Active learning on graph hierarchies. We now test the effectiveness of our proposed active learning labeling on graph hierarchies, and show a comprehensive comparison with several baselines in Fig. 5. We directly measure the accuracy of the output labels given an increasing percentage of manual annotations selected by different acquisition functions on MOT17. The vertical axis shows the increase in HOTA over the baseline performance, i.e., the performance without active learning. Oracle is our upper bound and shows the performance if all possible labels are set to their correct value by the annotator. Every method has two different labeling schemes: one for detections (nodes), and one for associations (edges) as explained in Sec. 3.3. In addition to our method, we explore active learning strategies commonly used in object detection, including Coreset [57], Ensemble [5], and Consistency over augmented inputs [21]. The naming is based on the active learning strategy for detection in the first part and the strategy for association in the second part. For instance, EntImg|Rand. indicates image level entropy-based selection for detection and random selection for association.
The methods shown in red shades and circle markers compare three detection labeling methods while keeping association label selection random. Box-level labeling is far superior to commonly-used image-level labeling, or random labeling. This highlights the potential of performing active learning at the object level rather than the frame level. The methods in blue shades and triangle markers all use our proposed active learning selection on graph hierarchies for the association, and compare several detection selection schemes. The Coreset [57], which prioritizes diversity of the data over uncertainty, emerges as the poorest performing approach. This observation underlines that, when accurate predictions are available, an uncertainty-based approach proves to be a more effective complement to pseudo-labels. Box-level entropy, ensembles, and consistency all perform similarly. Finally, our method, with both detection and association label selection on graph hierarchies is superior to all compared baselines, staying at only 1.2 points from Oracle performance. Notably, we also unify the AL strategy for detection and association with our approach.
@@ 4.6 Putting Everything Together: SPAM Labels in Action
Model Labels Pseudo-labels Active Learn. Ann. Effort $\downarrow$ HOTA $\uparrow$ MOTA $\uparrow$ IDF1 $\uparrow$
@@
Training trackers with our labels. In this experiment, we ablate the effects of our SPAM pipeline on the actual usecase of training a model with our provided labels. We ablate label quality under the following settings: (i) Synthetic training, (ii) Retraining our models on our Pseudo-labels, (iii) Active learning based acquisiton and annotation of a small percentage of real data. We train state-of-the-art online tracker ByteTrack [77] from scratch with the aforementioned sets of labels and report its performance on the MOT17 validation set in  Tab. 4. The goal performance we want to reach is the one achieved by the model trained on groundtruth annotations (row 1). Row 2 shows our model trained only on synthetic data, which is 3 HOTA points below our goal. Using pseudo-label based training of our model increases performance by almost a HOTA point, and the ground truth level performance is achieved by also using our active learning method to select which samples to label. By annotating only $3.3\%$ of the data, we can achieve the same results as the model trained on the ground truth data with $100\%$ annotation effort. This shows the potential of our SPAM in reducing annotation effort for multi-object tracking datasets.
@@
How much data do we really need to label? In this final experiment, our goal is to assess the impact of our entire pipeline by answering the question: How much data do we actually need to label in order to achieve the same performance as a model trained on the entire groundtruth training dataset? For that, we choose two well-established trackers, ByteTrack [77] and GHOST [56], and train them from scratch using our SPAM labels. We report their performance on MOT17, MOT20, and DanceTrack test sets in  Tab. 5. We first confirm that our results on the validation set also generalize to the test set, and that we can reach GT performance with $3.3\%$ of annotation effort on MOT17. Note that for MOT17 our annotations yield a slight improvement over GT, which can be attributed to observed annotation noise in the original GT, particularly in moving sequences (further details can be found in the supplementary material). MOT20 and DanceTrack are known to be more challenging datasets [59], which is confirmed by the fact that the trackers still miss a few points when trained on only $3.3\%$ of manually annotated data. Given that we cannot do a full analysis of different percentages because we are working on the test sets, we choose two percentages of manually annotated data on which to train our trackers: $10\%$ for MOT20, and $20\%$ for DanceTrack. In both cases, we achieve on-par performance with the trackers trained on full groundtruth data. This clearly shows the potential in properly leveraging synthetic data, pseudo-labels from a strong model, and active learning on graph hierarchies in order to label multi-object tracking data, where we can save $96.7\%$ of the annotation effort on MOT17, $90\%$ of the annotation effort on MOT20 and $80\%$ of the annotation effort on DanceTrack.
@@ 5 Conclusion
We have introduced SPAM, a labeling engine to efficiently obtain trajectory annotations from videos. We have shown that SPAM is able to produce annotations that yield state-of-the-art tracking performance with a significantly reduced human annotation effort. Through our extensive experiments we have shown that (i) our synthetic pretraining + pseudo-label finetuning recipe yields a strong, annotation-free, labeler; (ii) our combination of active learning + pseudo-labeling for annotation generation yields high quality and efficient annotations; and (iii) that graphs are an excellent tool to provide comprehensive modelling for trajectory annotation. We believe that SPAM paves the way to make large-scale cheap annotation of tracking datasets a reality, which will in turn power the next generation of data-hungry tracking algorithms.
@@
Appendix 0.A Additional Details about SPAM
0.A.1 GNN Model
Our hierarchical graph-neural network (GNN) utilizes the time-aware message passing framework introduced in [9]. We use a hierarchy of 10 GNNs operating over a video clip of 512 frames. At the first level, we employ a $GNN_{node}$, which considers a temporal window of 10 frames for each node. Subsequent levels, denoted as $GNN_{edge}$, cover 2, 4, …, up to 512 frames. We follow a simple and unified approach for all hierarchy levels: $GNN_{node}$ and $GNN_{edge}$ levels share the same input features and network architecture (as presented in Sec 3.3 of the main paper). Moreover, $GNN_{edge}$ levels also share learnable weights across hierarchy levels. Overall, we rely on a lightweight GNN hierarchy with only approximately 65K parameters running at 20 frames per second (FPS).
0.A.2 Active Hierarchical Labeling
0.A.3 Implementation Details
Experimental setup. After obtaining labels with SPAM, we train two well-established online trackers ByteTrack [77] and GHOST [56] to evaluate our label quality in the main paper. Our main goal is to compare our label quality with ground truth label quality, therefore, we start trainings from scratch and only train on the dataset of interest (200 epochs for MOT17, 150 epochs for MOT20 and 100 epochs for DanceTrack). During testing, we follow their default parameters [77, 56].
Appendix 0.B Additional Experiments
@@
0.B.1 Frame-based Labeling and FPS
@@
0.B.2 Label Noise for Moving Cameras
In our experiments detailed in Sec 4.6 of the main paper, we found that on MOT17, our annotations with a budget of 3.3% yield slightly better performance compared to the ground truth. We attribute this observation to annotation noise in the original ground truth, particularly in sequences with moving cameras. To illustrate this, we provide detailed results on MOT17 for each sequence with ByteTrack and GHOST in Fig. 7. We compare relative HOTA scores between trainings on ground truth and SPAM labels. Positive scores indicate superior performance with SPAM labels, while negative scores indicate that ground truth labels result in better performance. Overall, with a 3.3% budget, ground truth exhibits better performance on static camera sequences, while SPAM outperforms ground truth labels on moving camera sequences. We speculate that this phenomenon could be due to the additional noise in ground truth labels on moving sequences, which may be introduced by the interpolation-based labeling strategy reported in [18].
0.B.3 Alternative Evaluation Setting
As demonstrated in the literature on pseudo-labeling, active learning, and noisy labels, labels do not contribute equally when training models. Therefore, our main goal is not to achieve perfect labels but to label data with minimal annotation effort while enabling optimal tracking performance. Thus, we train SOTA trackers with our labels and compare the results with ground truth training in our experiments similar to an AL evaluation setting. For completeness, we also report tracking scores reached by our labels directly as an alternative setting here. We obtain 84.1 MOTA, 72.2 HOTA, and 90.1 IDF1 with our labels (3.3% annotation effort) on MOT17. This shows that the labels are indeed not completely covering groundtruth. However, results from the Table 5 of the main paper show that training on those labels gives the same results as training on full groundtruth. Thus, even though the labels are not covering the full ground-truth, allocating more annotation resources will not improve the training performance due to diminishing returns. This demonstrates that perfect labels are not necessary to reach optimal tracking performance and our experimental setting is well-suited for our goal.